# How Do Multilingual Models Remember? Investigating Multilingual Factual Recall Mechanisms

### v1

@@ Abstract
Large Language Models (LLMs) store and retrieve vast amounts of factual knowledge acquired during pre-training. Prior research has localized and identified mechanisms behind knowledge recall; however, it has primarily focused on English monolingual models. The question of how these processes generalize to other languages and multilingual LLMs remains unexplored. In this paper, we address this gap by conducting a comprehensive analysis of two highly multilingual LLMs. We assess the extent to which previously identified components and mechanisms of factual recall in English apply to a multilingual context. Then, we examine when language plays a role in the recall process, uncovering evidence of language-independent and language-dependent mechanisms.
How Do Multilingual Models Remember?
Investigating Multilingual Factual Recall Mechanisms
Constanza Fierro   Negar Foroutan   Desmond Elliott   Anders Søgaard University of Copenhagen    EPFL
@@ 1 Introduction
Large Language Models (LLMs) acquire a vast amount of factual knowledge during pre-training, including many propositional facts of the form “The capital of France is __” Petroni et al. (2019). Such acquisition also occurs in multilingual language models (Kassner et al., 2021; Jiang et al., 2020; Yin et al., 2022), although with considerable variation in performance depending on the input language. This raises an important question: Is this variation caused by the knowledge being stored in a language-specific manner, or is it related to the exact phrasing sensitivity? (Elazar et al., 2021) If trust turns on knowledge recall (Grasswick, 2010; Hawley, 2012; Nguyen, 2022), such assessments will be integral to establishing whether multilingual models are trustworthy.
Mechanistic interpretability aims to understand how and why models process information and make predictions at an internal structural level (Elhage et al., 2022; Nanda et al., 2023; Olah et al., 2020). Recent works have focused on identifying the specific model components responsible for storing acquired knowledge (Meng et al., 2022; Sharma et al., 2024), and understanding how this knowledge is retrieved internally during inference (Geva et al., 2023; Chughtai et al., 2024). However, so far, this research has been exclusively in English and largely limited to autoregressive LMs, except for Sharma et al. (2024), who study these mechanisms in Mamba (Gu and Dao, 2023), a state-space LM.
The acquisition of factual knowledge is usually studied by assessing whether a model predicts the expected fact, given a subject and relation as input, e.g. “[Y], which is the capital of France”. This setup was originally proposed for masked LMs Petroni et al. (2019), where the object to be predicted could be positioned anywhere within the input. With the increasing popularity of decoder-only models (Brown et al., 2020), however, the input format has shifted. For these models, the placeholder (e.g., “[Y]”) has to be placed at the end of the sentence, so that the model predicts the next token, which is the object (Meng et al., 2022; Yu et al., 2023; Geva et al., 2023; Hase et al., 2023; Wang et al., 2024; Cohen et al., 2024). Clearly, this approach presents challenges for languages that do not follow a subject-verb-object (SVO) word order. In languages like Turkish, Japanese, Korean, and Persian, which predominantly use a subject-object-verb (SOV) structure, it can be unnatural and difficult to craft inputs that fit this template for studying acquired factual knowledge.
In this paper, we study the mechanisms of factual recall in multilingual LLMs, focusing on two architectures: the decoder-only model XGLM and the encoder-decoder model mT5; and centering on four key questions: (1) Are the factual recall mechanisms and localization found for English LMs also present in multilingual LLMs? (2) Do these mechanisms function the same way for different languages? (3) Does the model architecture—encoder-decoder or decoder-only—affect these mechanisms? (4) When language is encoded, are there language-dependent mechanisms at play? To address (1)–(4), we first conduct a causal tracing analysis (§3) to determine whether the early MLP modules processing the final subject token are as decisive in multilingual models as in English-centric models (Meng et al., 2022). In XGLM, this effect is absent, while in mT5, we observe a strong causal impact from both the first and last subject tokens across all layers of the encoder.
@@
Second, we analyze the mechanisms described in Geva et al. (2023), with a focus on when information from the subject and relation flows to the last token, and when the object is extracted from the subject representation (§4). We find that information flows similarly in monolingual and multilingual decoder-only models, while for mT5, there are discrepancies. In terms of the extraction event, in XGLM, both feed-forward and attention sublayers contribute to it, in contrast to English autoregressive models, where attention modules do most work. In mT5, this mechanism is primarily performed by the cross-attention, occurring later in the process and in a more localized manner compared to decoder-only models.
Finally, we use activation patching (Zhang and Nanda, 2024; Ghandeharioun et al., 2024) and insert the last token representation from an English input into the forward pass of another language (§5); this way, we are able to characterize more fully the multilingual recalling process. We find that in XGLM there are two distinct phases. First the relation representation propagates to the last token, and then the output language. While for mT5, both the relation and language propagate in the later layers, with the relation being encoded slightly earlier. Additionally, we show that the subject and relation representations are language-independent, as they can be interchanged and used in different languages; while the extraction event is language-specific. The findings of this work contribute to the understanding of crosslingual factual knowledge transfer and consistency, and can further inform knowledge editing methodologies, and trustworthiness assessments in multilingual models.
@@ 2 Experimental Setup
We focus on the basic form of factual knowledge recall, where LMs are tasked with predicting the correct object $o$ for a given subject $s$ and a relation $r$. These $(s,r,o)$ triplets are obtained from WikiData, and natural language templates are used to describe the relation, with placeholders for the subject and object. For the relation born-in, a template could, e.g., be “[X] was born in [Y]”, where [X] represents the placeholder for the subject and [Y] is the object to be predicted by the model.
We use the mParaRel dataset (Fierro and Søgaard, 2022), with triplets and templates for 45 languages. [^1] We augment mParaRel by fetching WikiData aliases of the objects to be able to identify more cases of correct object prediction. We filter out trivial examples where the object is directly mentioned in the query, e.g., South Asia belongs to the continent of. For our analysis, we select 10 typologically diverse languages representing a variety of scripts, language families, and word orders, including SVO, SOV, and VSO structures: English (en), Spanish (es), Vietnamese (vi), Turkish (tr), Russian (ru), Ukrainian (uk), Japanese (ja), Korean (ko), Hebrew (he), Persian (fa), and Arabic (ar).
We analyze two different architectures of multilingual language models, XGLM (Lin et al., 2021),
Language XGLM mT5
en 1812 1543
es 1380 1192
vi 1646 993
tr 418 1058
ru 830 683
uk 213 456
ko 116 630
ja 6 358
he 13 565
fa 7 406
ar 811 488
Table 1: Number of memorized knowledge triplets $(s,r,o)$.
a decoder-only architecture with 7.5B parameters and 32 layers, and mT5-xl (Xue et al., 2021), an encoder-decoder architecture of 3.7B parameters and 24/24 encoder/decoder layers. For XGLM, we only use the templates in mParaRel that have the object placeholder at the end of the sentence, while for mT5, we use all the templates available. XGLM pre-training data includes 30 languages but it does not include uk, he, and fa; while mT5 includes 101 languages and covers all the ones studied here.
@@
As we aim to investigate the process of knowledge recall, and similarly to previous work (Meng et al., 2022; Geva et al., 2023), we only consider examples where the model predicts the correct object completion. Since mParaRel provides multiple templates (paraphrases) for each relation, we greedy-decode for every available template corresponding to a given triplet and check for an exact match. [^2] We perform greedy decoding of a maximum of 50 tokens and check for exact match of any of the object aliases. In practice, the 50 tokens generation is an overestimate, as on average, around 0.5 tokens are generated prior to the target object (for some templates an article or other filling tokens are generated before predicting the object). If multiple templates trigger the correct object, we randomly select one for the analyses. Once an exact match is found, any extra tokens generated before the object are added to the example input text. This ensures that when we feed again the example to the model, the next token predicted is the first token of the object. [^3] For mT5, we feed the mParaRel input into the encoder with a sentinel token in the object placeholder. In the decoder, we provide the beginning-of-sequence token followed by the sentinel token. We only check the tokens generated next to the first sentinel token, as the pre-training task of the model is to generate the text for each sentinel in the input, the decoder usually continues to generate answers for other sentinel tokens. Any tokens generated before the target object, are added to the decoder tokens since adding these to the encoder does not ensure that the next token predicted will be the object. Table 1 presents the number of examples for which the correct object is predicted. Due to the limited number of examples memorized by XGLM in ja, he, and fa, we exclude these languages from our main analysis of XGLM.
@@ 3 Causal Tracing
We first analyze which hidden states in the model’s computation are more important than others when recalling a fact. Following Meng et al. (2022), we trace the causal effects of hidden states using causal mediation analysis (Pearl, 2022). To do this, we feed the model a corrupted input and selectively restore a particular component in the computational graph, such as the attention output of layer $l$. We then measure how much of the original probability (from the uncorrupted input) is recovered, thereby assessing the indirect influence of the restored component on the model’s prediction. Specifically, we corrupt the input by adding Gaussian noise to the subject tokens, [^4] We follow Meng et al. (2022) and add $\epsilon\sim\mathcal{N}(0,(3\sigma)^{2})$, with $\sigma$ being the standard deviation of the subjects tokens embeddings from the data used. and observe the probability $\hat{\mathbb{P}}$ of the originally predicted token. Then, we run inference again on the corrupted input, but this time, we restore a specific hidden state in the model and track the probability ^$\hat{\mathbb{P}}_{\text{restored}}$ of the originally predicted token. We study the indirect effect $\text{IE}=\hat{\mathbb{P}}_{\text{restored}}-\hat{\mathbb{P}}$.
Consider a transformer model with $L$ layers; let $h_{t}^{l}$ be the representation of the token $t$ at layer $l$. When the model is an encoder-decoder, let $e_{i}^{l}$ be the representation of the encoder layer $l$ for the $i$-th token in the input. Then, the encoder layer computes $h_{t}^{l+1}=h_{t}^{l}+s^{l}+f^{l}$ and the decoder $h_{t}^{l+1}=h_{t}^{l}+s^{l}+c^{l}+f^{l}$, where $s^{l}=\text{Self Attn.}(h_{0}^{l}\ldots h_{t}^{l})$, $c^{l}=\text{Cross Attn.}(h_{t}^{l},e_{0}^{l}\ldots e_{n}^{l})$ and $f^{l}=\text{MLP}(h_{t}^{l}+s^{l}+c^{l})$. If decoder-only, then $c^{l}$ does not apply.
Let $h_{t}^{l}$ be the hidden state from the clean run and $\overline{h}_{t}^{l}$ that of the corrupted run. We restore a state by setting $\overline{h}_{t}^{l}\leftarrow h_{t}^{l}$, and similarly, we restore the self-attention layers contribution by setting $\overline{s}^{l}\leftarrow s^{l}$ for all the attention modules in a window of size $w$ (analogous for $c$ and $f$). [^5] We use windows because, generally, the contributions of the sub-layers are gradual (Geva et al., 2021). In other words, multiple layers contribute the same behavior to the residual, and their sum produces the observed effect. Following Meng et al. (2022), we set $w=10$ for XGLM, and $w=6$ for mT5 (as it has 24 layers). We repeat the corrupted run with restoration ten times, each with different noise samples, and report the average.
@@
We compare our results in multilingual LLMs to those of Meng et al. (2022), who analyzed the factual recall of GPT-2 XL in English, [^6] They used the CounterFact dataset, which is based on the ParaRel dataset. Thus, the English data is the same as the one we use here. and reached two key conclusions. First, they identified an “early site”, where the MLPs processing the last subject token in the early and middle layers play a crucial role in recovering from input corruption. Second, they found a “late site”, where the attention modules processing the last token in the later layers also significantly contribute to prediction recovery. [^7] While they consider the late site unsurprising, as it directly precedes the final prediction, this observation primarily applies to hidden state restoration, not necessarily attention restoration. We, however, interpret the late-site attention as aligning with what Geva et al. (2023) referred to as the extraction event (§4).
We present the causal analysis plots for each model and language in Appendix B and present our main observations here. We do not find for XGLM an early site of MLP modules nor for English or any of the other languages. Thus, we conclude that this feature is specific to English-centric models, as Sharma et al. (2024) also observed it for another English-centric architecture (Mamba). We do observe a causal effect of the last subject token when restoring the state in the early layers. This is somewhat expected as it reverts the corruption performed in the input, but it is important to note until which layer it is possible to recover from this corruption only by restoring the last subject token. It also highlights the relevance of the last subject token in decoder-only models for all languages.
@@
Figure 1: Attention knockout between the last token and a given set of tokens. Each layer represents the effect of the knockout on a window of $w$ layers. Top XGLM ($w=6$), bottom mT5 ($w=4$).
For mT5, we do observe a causal effect from MLP modules of the subject tokens, but it is not an early site as in GPT. Rather it is a causal effect along all the layers of the encoder, in both the first and last token of the subject, with some languages having a stronger effect in the last subject token. In terms of late site, we find it in the self attention for XGLM, and in the cross attention layers in mT5. For both mT5 and XGLM, we also observe a late site of the MLP at the last token, though this effect is smaller than that of the attention or the subject.
We conclude that the multilingual models behave in a similar way for the different languages, finding that the causal effects of both MLP and attention modules are positioned the same, and that for decoder-models the last subject token (at the early layers) is critical for the recalling, while for encoder-decoder models the MLPs processing both the first and last subject token encode the necessary information (at any encoder layer).

### v2

# How Do Multilingual Language Models Remember Facts?

@@ Abstract
Large Language Models (LLMs) store and retrieve vast amounts of factual knowledge acquired during pre-training. Prior research has localized and identified mechanisms behind knowledge recall; however, it has only focused on English monolingual models. The question of how these mechanisms generalize to non-English languages and multilingual LLMs remains unexplored. In this paper, we address this gap by conducting a comprehensive analysis of three multilingual LLMs. First, we show that previously identified recall mechanisms in English largely apply to multilingual contexts, with nuances based on language and architecture. Next, through patching intermediate representations, we localize the role of language during recall, finding that subject enrichment is language-independent, while object extraction is language-dependent. Additionally, we discover that the last token representation acts as a Function Vector (FV), encoding both the language of the query and the content to be extracted from the subject. Furthermore, in decoder-only LLMs, FVs compose these two pieces of information in two separate stages. These insights reveal unique mechanisms in multilingual LLMs for recalling information, highlighting the need for new methodologies—such as knowledge evaluation, fact editing, and knowledge acquisition—that are specifically tailored for multilingual LLMs.
How Do Multilingual Language Models Remember Facts?
Constanza Fierro∗†   Negar Foroutan   Desmond Elliott   Anders Søgaard University of Copenhagen    EPFL
**footnotetext: Correspondance: Constanza Fierro <c.fierro@di.ku.dk>.
@@ 1 Introduction
Large Language Models (LLMs) learn extensive factual knowledge during pre-training, including propositional facts like “The capital of France is __” Petroni et al. (2019). While multilingual models also acquire such knowledge, their performance varies significantly across languages (Kassner et al., 2021; Jiang et al., 2020; Yin et al., 2022), raising questions about whether this variation stems from language-specific storage or phrasing sensitivity (Elazar et al., 2021). Such assessments will be critical for determining the trustworthiness of multilingual LLMs, if trust relies on knowledge (Grasswick, 2010; Hawley, 2012; Nguyen, 2022).
Mechanistic interpretability research has begun uncovering how models store and retrieve knowledge internally, with recent studies identifying specific components for knowledge storage (Meng et al., 2022; Sharma et al., 2024) and retrieval mechanisms (Geva et al., 2023; Chughtai et al., 2024). However, they have focused exclusively on English LLMs, mainly autoregressive ones. [^1] Except Sharma et al. (2024), who extend these analyses to Mamba (Gu and Dao, 2023), a state-space language model. Encoder-decoder architectures, which could enable better cross-lingual representations (Li et al., 2024), remain underexplored. Additionally, recent studies have shown that LLMs share circuits across languages for specific tasks (Ferrando and Costa-jussà, 2024; Zhang et al., 2025), but these are limited to syntactic tasks and do not address how concepts are represented or recalled cross-lingually. While Dumas et al. (2024) examined how disentangled language is from concepts using a translation task, here we analyze fact recall, which allows us to offer broader insights into how language is encoded.
In this paper, we study the mechanisms of factual recall in multilingual LLMs, focusing on two architectures: decoder-only (XGLM and EuroLLM) and encoder-decoder (mT5). We analyze a simple form of information extraction, where the input contains a subject and a relation, and the model predicts the corresponding object (e.g. “Paris” in the earlier example). Our analysis centers on three key questions: (1) Does the localization of factual knowledge in English LLMs extend to multilingual LLMs? (2) Are the factual recall mechanisms found for English LLMs also present in multilingual LLMs? (3) When does language play a role in the recall mechanism?
To address the first question, we use causal tracing analysis to assess if early MLP modules processing the final subject token are as decisive in multilingual models as in English-centric models (Meng et al., 2022). Our results (§3) show that EuroLLM and mT5 exhibit strong causal effects for all subject tokens, with EuroLLM exhibiting this in earlier layers and mT5 across all encoder layers. In XGLM, the last subject token has a stronger effect, but with two early sites, unlike English GPT’s single site. Additionally, in all models, MLPs in later layers are decisive in recovering factual information, a contrast to Meng et al. (2022) findings.
@@
Next, we investigate the second question by analyzing the mechanisms described in Geva et al. (2023); namely, the three-step process, where the relation information flows to the last token, then the subject flows, and finally the attention layers extract the object (§4). We find that in multilingual decoder-only models the information flows similarly to monolingual models, but differently for the encoder-decoder model mT5. In terms of the extraction event, in decoder-only LLMs, both feed-forward and attention sublayers contribute to it, unlike English autoregressive models, where attention modules dominate. In mT5, this mechanism is primarily performed by the cross-attention. Overall, our findings indicate that some of the localization (§3) and mechanisms (§4) of fact retrieval in English LLMs generalize to multilingual LLMs, but with key variations.
Finally, to address the third question, we investigate where language plays a role within the three-step process described by Geva et al. (2023), to characterize how and where factual knowledge cross-lingual transfer may occur. Using activation patching (Zhang and Nanda, 2024; Ghandeharioun et al., 2024) we insert the intermediate representation of the last token from an English forward pass into the forward pass of another language (§5). Our results reveal that the last token acts as a Function Vector (FV) Todd et al. (2024), encoding both the relation and the output language, which is then applied to the subject in the context. Crucially, the fact that the FV formed in one language can be used with an input in another language demonstrates that the subject and relation representations are largely language-independent, while the extraction event is language-specific. Furthermore, in decoder-only LLMs (XGLM, EuroLLM), the FV is constructed in two distinct phases: first, it encodes only the relation, and later, the language is incorporated.
These findings advance our understanding of factual recall, with implications for cross-lingual knowledge transfer, knowledge editing, and trustworthiness in multilingual LLMs. Our results on language flow open new avenues for studying whether models ‘think’ in English (Wendler et al., 2024), by examining how the FV is altered across languages. Additionally, the late-site causal effect of MLPs and their dominance in the extraction phase suggest that fact editing techniques and knowledge evaluations must extend beyond early MLPs and attention layers (Tamayo et al., 2024).
@@ 2 Experimental Setup
We focus on a simple form of factual knowledge recall, where LMs are tasked with predicting the correct object $o$ for a given subject $s$ and a relation $r$. These $(s,r,o)$ triplets are obtained from WikiData, and natural language templates are used to describe the relation, with placeholders for the subject and object.222 For example, the relation born-in could use the template “[X] was born in [Y]”, where [X] is the subject and [Y] is the object to be predicted. For our analysis, we select 10 typologically diverse languages representing various scripts, families, and word orders: English (en), Spanish (es), Vietnamese (vi), Turkish (tr), Russian (ru), Ukrainian (uk), Japanese (ja), Korean (ko), Hebrew (he), Persian (fa), and Arabic (ar). See Table 2 for language characteristics.
Models
We analyze decoder-only and encoder-decoder architectures. [^3] For decoder-only models, we only use the templates in mParaRel that have the object placeholder at the end of the sentence, while for encoder-decoder, we use all the templates. The decoder-only LLMs are XGLM (Lin et al., 2021),
XGLM EuroLLM mT5
en 1812 2332 1543
es 1380 1913 1192
vi 1646 779 993
tr 418 799 1058
ru 830 1680 683
uk 213 1244 456
ko 116 308 630
ja 6 42 358
he 13 107 565
fa 7 31 406
ar 811 1790 488
Table 1: Number of facts $(s,r,o)$ correctly predicted.
with 7.5B parameters and 32 layers, and EuroLLM (Martins et al., 2024) with 9B parameters and 42 layers; the encoder-decoder mT5-xl (Xue et al., 2021) has with 3.7B parameters and 24 encoder-decoder layers. The pre-training data varies across models: mT5 is pretrained on 101 languages, covering all the languages in our study; XGLM covers 30 languages, excluding uk, he, and fa; while EuroLLM covers 35 languages, excluding vi, he, and fa.
Data
We use the mParaRel dataset (Fierro and Søgaard, 2022), which includes triplets and templates for 45 languages. These templates are machine translations of the ParaRel English templates (Elazar et al., 2021). [^4] We augment the objects in mParaRel, filter out trivial examples, and when there are enough examples, we use a crosslingual subset (see Appendix B).
@@
To investigate the process of knowledge recall we only consider examples where the model predicts the correct object completion (Meng et al., 2022; Geva et al., 2023). Since mParaRel provides multiple paraphrased templates for each relation, we greedy-decode for every available template corresponding to a given triplet and check for an exact match. If multiple templates yield a match, we randomly select one for the analysis. In cases where an article or other filler tokens precede the object, we include these tokens in the input text to ensure that when the example is fed into the model for our analysis, the next predicted token is the first token of the object (Implementation details in Appendix B.1). Table 1 presents the number of examples for which the correct object is predicted. We exclude languages with too few examples from our analysis, namely ko, ja, he, and fa for XGLM, and ja, he, and fa for EuroLLM.
Notation
Given a transformer model with $L$ layers, let $h_{t}^{l}$ be the representation of the token $t$ at layer $l$. When the model is an encoder-decoder, let $e_{i}^{l}$ be the representation of the encoder layer $l$ for the $i$-th token in the encoder input. Then, the encoder layer computes $h_{t}^{l+1}=h_{t}^{l}+s^{l}+f^{l}$ and the decoder $h_{t}^{l+1}=h_{t}^{l}+s^{l}+c^{l}+f^{l}$, where $s^{l}=\text{Self Attn.}(h_{0}^{l}\ldots h_{t}^{l})$, $c^{l}=\text{Cross Attn.}(h_{t}^{l},e_{0}^{l}\ldots e_{n}^{l})$ and $f^{l}=\text{MLP}(h_{t}^{l}+s^{l}+c^{l})$. If decoder-only, then $c^{l}$ does not apply.
@@ 3 Causal Tracing
We first analyze which hidden states in the model’s computation are more important than others when recalling a fact. Following Meng et al. (2022), we trace the causal effects of hidden states using causal mediation analysis (Pearl, 2022). Let $\mathbb{P}(o)$ be the probability of the predicted object token, and $LS(o)$ its logit score. We corrupt the input by adding Gaussian noise to the subject tokens, [^5] We follow Meng et al. (2022) and add $\epsilon\sim\mathcal{N}(0,(3\sigma)^{2})$, with $\sigma$ being the standard deviation of the subjects tokens embeddings from the data used. and observe the corrupted probability ~$\tilde{\mathbb{P}}(o)$ of the originally predicted token. Then, we run inference again on the corrupted input, but this time, we restore a specific hidden state in the model and track the probability ~$\tilde{\mathbb{P}}_{\text{restored}}(o)$. We study the indirect effect of such component as $\text{IE}_{\mathbb{P}}=\tilde{\mathbb{P}}_{\text{restored}}(o)-\tilde{\mathbb{P}}(o)$, or if using logits $\text{IE}_{LS}=\tilde{LS}(o)-\tilde{LS}_{\text{restored}}(o)$. Specifically, we restore a state by setting $\tilde{h}_{t}^{l}\leftarrow h_{t}^{l}$, where $h_{t}^{l}$ is the hidden state from the clean run and $\tilde{h}_{t}^{l}$ that of the corrupted run; and similarly, we restore the self-attention layers contribution by setting $\tilde{s}^{l}\leftarrow s^{l}$ for all the attention modules in a window of size $w$ (analogous for $c$ and $f$). [^6] We use windows because, generally, the contributions of the sub-layers are gradual (Geva et al., 2021). In other words, multiple layers contribute the same behavior to the residual, and their sum produces the observed effect. We set $w$=$4$ for XGLM, $w$=$5$ for EuroLLM, and $w$=$3$ for mT5, to restore $\sim 12\%$ layers. We repeat the corrupted run with restoration ten times with different noise samples, and report the average.
@@
We compare our results in multilingual LLMs to those of Meng et al. (2022), who analyzed the factual recall of GPT-2 XL in English [^7] They used the CounterFact dataset, which is based on ParaRel. Thus, their English data is the same as our data. and reached two key conclusions: (1) they identified an “early site”, where the MLPs processing the last subject token in the early and middle layers play a crucial role in recovering from input corruption; and (2), they found a “late site”, where the attention modules processing the last token in the later layers also significantly contribute to prediction recovery. [^8] While they consider the late site unsurprising, as it directly precedes the final prediction, this observation primarily applies to hidden state restoration, not necessarily attention restoration. We, however, interpret the late-site attention as aligning with what Geva et al. (2023) referred to as the extraction event (§4).
@@
Figure 1: Attention knockout between the last token and a given set of tokens. Each layer represents the effect of the knockout on a window of $w$ layers. Top EuroLLM ($w=7$), bottom mT5 ($w=4$). XGLM in Figure 21.
We present the causal analysis plots for each model and language in Appendix C and present our main observations here. [^9] In line with Zhang and Nanda (2024) we find that analyzing the $\text{IE}_{\mathbb{P}}$ overestimates the causal effect of some tokens over others. For example, in XGLM by looking at the $\text{IE}_{\mathbb{P}}$ we would not find an early site, which is present when studying the $\text{IE}_{LS}$ (see Figure 11 vs Figure 9), or in EuroLLM the last subject token would seem more relevant than the other subject tokens (see Figure 15 vs Figure 13). So we base our main observations using $\text{IE}_{LS}$. Our results indicate that the last subject token does not consistently have a stronger effect on information recovery in early MLPs. In EuroLLM and mT5, all subject tokens exhibit similar effects, while in XGLM, the last subject token has a stronger influence. In EuroLLM, the early site spans half of the initial layers, whereas in mT5, it covers all encoder layers. For XGLM, the early site occurs twice: once in the initial layers and again in the middle layers. Regarding the late site, we observe a strong causal effect in MLP layers before the very last layers in all models. The layers where these sites occur are consistent across languages, with only a slight variation in the late site’s endpoint for EuroLLM.
These findings suggest that some, but not all, conclusions from English LLMs generalize to multilingual LLMs. The MLP early site extends across models, and the MLP late site, which was not observed by Meng et al. (2022), was present in all models studied. This has implications for knowledge localization and fact editing methodologies, which should account for both early and late MLPs in multilingual contexts.